OpenStudy (anonymous):

how do I find out what 10 1/2% of 60?

OpenStudy (mertsj):

Do you know that, in mathematics, "of" means times?

OpenStudy (anonymous):

yes I do

OpenStudy (jdoe0001):

\(\bf {\color{red}{ a}}\%\ of \ b=\cfrac{{\color{red}{ a}}}{100}\times b\qquad \qquad {\color{red}{ 10\frac{1}{2}}}\%\ of \ 60=\cfrac{{\color{red}{ 10\frac{1}{2}}}}{100}\times 60\)

OpenStudy (anonymous):

0.105*60=6.3

OpenStudy (anonymous):

is 6.3 the correct answer?